# Towards Learning Stochastic Population Models by Gradient Descent

### v1

@@ 5.1 Reparametrization
Optimizing in this space means that a step in $\mathbf{r}$ between, e.g., $0.1$ and $0.2$ is the same as between $1$ and $2$. The specific shifting and scaling ensure (1) that the value $r_{i}=0$ is mapped to 00 and (2) that the values in a sensible range (around $10^{-4}$ to $10^{2}$) are sufficiently spread. This way, the sensitivity of the response wrt. changes in $\mathbf{r}$ is decreased, aiding the optimization (cf. Figure 1). Specifically, in the case of our stochastic gradient estimator, this allows us to set a single smoothing factor $\sigma$ for all dimensions, which would otherwise lead to an overly smoothed objective and occlude narrow minima.
5.2 Evaluation Setup
To identify the challenges and opportunities of gradient descent in the context of a stochastic simulation-based model inference, we evaluate the convergence of our four problem formulations on recovering the SIR model as parametrized in Section 2. Our time-series reference data is generated by simulating the model until $t=1$ and collecting state snapshots at $100$ discrete, equidistant simulation times (although we generally require neither equidistance nor completeness). For optimization, we employ the stochastic gradient estimator introduced in Section 3 and combine it with the Adam gradient descent optimizer [14]. For each problem, we manually determined hyperparameters (sample size $n$, smoothing factor $\sigma$, and learning rate $\eta$) that achieved good results. In the order of the problems from Section 5, these are $(100,0.2,1)$, $(1000,1,1)$, $(100,0.2,0.1)$, and $(100,0.2,0.5)$. Initial parameters are drawn from problem-specific uniform distributions. Our simple demonstration aims to minimize the root mean squared error (RMSE) between the reference and the simulation time-series, which is run for $20$ replications. Note that it is easily possible to change this objective, e.g., to minimizing Wasserstein distances on distribution estimates [28]. We repeat the optimization process $10$ times to account for the stochasticity.
6 Results and Discussion
The evaluation results are provided in Figure 2, which shows the mean convergence behavior over gradient descent steps on each problem, as well as the final model inferred by a chosen optimization run. For the Brute Force problem, the lowest RMSE of all structures is shown.
The Library of Reactions formulation yields a very precise fit to the input data but lacks parsimony. Convergence is attained fast, as the objective is smooth. Here, a parsimony-encouraging initialization, as with the horseshoe prior for Bayesian regression may prove beneficial [12], albeit introducing bias towards certain solutions.
@@ 6 Results and Discussion
Being completely smooth, the brute force Library of Systems approach is similar in convergence to the Library of Reactions. In contrast to the latter, it is able to recover the parsimonious original model. This indicates the ability of gradient descent to optimize a vast number of reaction systems at a time. Since the combinatorial explosion puts larger reaction systems out of reach, the main missing building block for this approach is a goal-driven exploration of structures.
Our initial results demonstrate a tradeoff between parsimony, goodness of fit and scalability. This is the result of different response surfaces and their amenability to gradient descent.
In all cases, the scaling of rate constants poses a problem, which can be dealt with by reparametrization (cf. Section 5.1). Whereas the rate constant space clearly places solutions of similar quality close to each other (cf. Figure 1), it is generally unclear which steps in the structure dimension (e.g., on the coefficients in $\mathbf{C}$) lead to lower loss. The simultaneous adjustment of both $\mathbf{C}$ and $\mathbf{r}$ further complicates solutions that try to (smoothly) enforce a certain model size. A major step towards better convergence would thus be a combined reparametrization of $\mathbf{C}$ and $\mathbf{r}$, which enables a goal-driven exploration of structures. Clearly, such a reparametrization must be approximate, and its existence is unclear, demanding further investigation. Promisingly, in the related case of learning (imperative) programs, first steps have been taken in this direction [19]. Besides parsimony, identifyability could be facilitated by constraining solutions on background knowledge, as for example derived from a conceptual model in a simulation study.
Beyond considering the challenges outlined above, future work may explore the application of other smooth gradient estimation schemes based on automatic differentiation, such as StochasticAD [2] or DiscoGrad [17]. Finally, the full potential of the simulation-based approach needs to be explored, e.g., by considering unmeasured variables and alternative loss functions.
Acknowledgements
JNK and AU acknowledge the funding of the Deutsche Forschungsgemeinschaft under Grant No.: $320435134$ (https://gepris.dfg.de/gepris/projekt/320435134); PA is supported by Grant No.: $497901036$ (https://gepris.dfg.de/gepris/projekt/497901036).

### v2

@@ 5.1. Reparametrization
Optimizing in this space means that a step in $\mathbf{r}$ between, e.g., $0.1$ and $0.2$ is the same as between $1$ and $2$. The specific shifting and scaling ensure (1) that the value $r_{i}=0$ is mapped to 00 and (2) that the values between $10^{-4}$ and $10^{2}$ are sufficiently spread. This way, the sensitivity of the response wrt. changes in $\mathbf{r}$ is decreased, aiding the optimization (cf. Figure 1). Specifically, in the case of our stochastic gradient estimator, this allows setting a single smoothing factor $\sigma$ for all dimensions, which would otherwise lead to oversmoothing and occlude narrow minima.
5.2. Evaluation Setup
To identify the challenges and opportunities of gradient descent in the context of a stochastic simulation-based model inference, we evaluate the convergence of our four problem formulations on recovering the SIR model as parametrized in Section 2. Our time-series reference data is generated by simulating the model until $t=1$ and collecting state snapshots at $100$ discrete, equidistant simulation times (although we generally require neither equidistance nor completeness). For optimization, we employ the stochastic gradient estimator introduced in Section 3 and combine it with the Adam gradient descent optimizer (Kingma and Ba, 2014). For each problem, we manually determined hyperparameters (sample size $n$, smoothing factor $\sigma$, and learning rate $\eta$) that achieved good results. In the order of the problems from Section 5, these are $(100,0.2,1)$, $(1000,1,1)$, $(100,0.2,0.1)$, and $(100,0.2,0.5)$. Initial parameters are drawn from problem-specific uniform distributions. Our simple demonstration aims to minimize the root mean squared error (RMSE) between the reference and the simulation mean time-series, the latter being determined from $20$ replications. Note that it is easily possible to change this objective, e.g., to minimizing Wasserstein distances on distribution estimates (Öcal et al., 2020). We repeat the optimization process $10$ times to account for the stochasticity.
6. Results and Discussion
The evaluation results in Figure 2 show the mean convergence behavior over gradient descent steps on each problem, as well as the final model inferred by a chosen optimization run. For the Brute Force problem, the lowest RMSE of all structures is shown.
The Library of Reactions formulation yields a very precise fit to the input data but lacks parsimony. Convergence is attained fast, as the objective is smooth. Here, a parsimony-encouraging initialization, such as the horseshoe prior for Bayesian regression may be beneficial (Jiang et al., 2022), albeit introducing bias towards certain solutions.
@@ 6. Results and Discussion
Being completely smooth, the brute force Library of Systems approach is similar in convergence to the Library of Reactions. In contrast to the latter, it is able to recover the parsimonious original model. This indicates the ability of gradient descent to optimize a vast number of reaction systems at a time. Since the combinatorial explosion puts larger systems out of reach, the main missing piece for this approach is a goal-driven exploration of structures.
Our initial results demonstrate a tradeoff between parsimony, goodness of fit, and scalability. This is the result of different response surfaces and their amenability to gradient descent. In all cases, the scaling of rate constants poses a problem, which can be dealt with by reparametrization (cf. Section 5.1). Whereas the rate constant space clearly places solutions of similar quality close to each other (cf. Figure 1), it is generally unclear which steps in the structure dimension (on the coefficients in $\mathbf{C}$) lead to lower loss. The simultaneous adjustment of both $\mathbf{C}$ and $\mathbf{r}$ further complicates solutions that try to (smoothly) enforce a certain model size. A major step towards better convergence would thus be a combined reparametrization of $\mathbf{C}$ and $\mathbf{r}$ that enables a goal-driven exploration of structures. Clearly, such a reparametrization must be approximate, and its existence is unclear, demanding further investigation. Promisingly, in the related case of learning (imperative) programs, first steps have been taken in this direction (Kusner et al., 2017). Besides parsimony, identifyability could be facilitated by constraining solutions on background knowledge, as for example derived from a conceptual model in a simulation study.
Beyond considering the challenges outlined above, future work may explore the application of other smooth gradient estimation schemes based on automatic differentiation, such as StochasticAD (Arya et al., 2022) or DiscoGrad (Kreikemeyer and Andelfinger, 2023). Finally, the full potential of the simulation-based approach needs to be explored, e.g., by considering unmeasured variables and alternative loss functions.